A 0.300-kg ball has a speed of 15.0 m/s. If its speed were doubled, what would be its kinetic energy

Select the correct one:

- the kinetic energy will be 67.5 joules

- The kinetic energy will be 135 joules

- The kinetic energy increases but it is not known to what extent

- The kinetic energy will be 33.75 joules
Answer #1

here,

as the kinetic energy is

kinetic energy = 0.50 *m * v^2

as v is doubles

the kinetic energy will be 2^2 = 4 times

kinetic energy = 0.50 * 0.30 * (15 * 2)^2

kinetic energy = 135 J

the kinetic energy is 135 J